OpenStudy (candra):

I've got two problems i'm stuck on...=-( -(6y+3)-(-5y-8)=3 is the first

OpenStudy (anonymous):

-6y-3+5y+8=3 -1y+5=3 5-3=1y 2=y

OpenStudy (anonymous):

this is where i was going with it in chat :p